OpenStudy (anonymous):

A) In the least-squares line y = 5 – 1x, what is the value of the slope? B) When x changes by 1 unit, by how much does y change?

OpenStudy (anonymous):

please explain..

OpenStudy (anonymous):

y=mx+b m= slope

OpenStudy (anonymous):

?

OpenStudy (anonymous):

The formula for a line is y=mx+b where m is the slope. Rewrite your equation like I have shown...

OpenStudy (anonymous):

so its y= 1x+5..?

OpenStudy (freckles):

what happen to the negative 1 in front of the x?

OpenStudy (anonymous):

it went bye bye

OpenStudy (anonymous):

???????

OpenStudy (anonymous):

y=-1x+5?

OpenStudy (anonymous):

better

OpenStudy (anonymous):

slope is that number sitting in front of the \(x\)

OpenStudy (anonymous):

so its -1?

OpenStudy (freckles):

keep in mind slope=rise/run rise effects the y number and run effects the x number

OpenStudy (anonymous):

ok... thanks !